OpenStudy (anonymous):

A proportional relationship between the number of pounds of cabbage (x) and the price in dollars (y) is graphed, and the ordered pair (5, 2) is on the graphed line. Part A: What is the price of 1 pound of cabbage? Show your work, including the proportion you used to determine the price. Part B: What does the ordered pair (10, 4) on the graph represent? Explain in words.

OpenStudy (phi):

when they say proportion, think ratios (i.e fractions) for example, the (5,2) means 5 pounds of cabbage and it cost 2 dollars write that as a fraction \[ \frac{ 5 \ pounds}{2 \ dollars}\] next , to answer Part A: What is the price of 1 pound of cabbage? we write another ratio 1 pound / y dollars (we don't know what y is

OpenStudy (phi):

set the ratios equal to each other \[ \frac{ 5 \ pounds}{2 \ dollars} = \frac{1 \ pound}{y} \] to solve for y, I would "flip" both sides , like this: \[ \frac{2 \ dollars}{ 5 \ pounds} = \frac{y} {1 \ pound} \]

OpenStudy (anonymous):

Ok i got that part @phi

OpenStudy (anonymous):

how do i do the rest @welshfella

OpenStudy (anonymous):

@Awolflover1